OpenStudy (anonymous):

Given the functions m(x) = 4x − 11 and n(x) = x − 10, solve m[n(x)] and select the correct answer below. m[n(x)] = 4x − 51 m[n(x)] = 4x − 29 m[n(x)] = 4x2− 51 m[n(x)] = 4x2 − 29

OpenStudy (anonymous):

Okay first off welcome to Open Study :)

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

No problem!

OpenStudy (anonymous):

Can you help me?

OpenStudy (anonymous):

I'll try!

OpenStudy (anonymous):

Okay so 4x*x would be what?

OpenStudy (anonymous):

4x^2

OpenStudy (anonymous):

and combine -10 and -11

OpenStudy (anonymous):

-21

OpenStudy (anonymous):

wait, in the choices are you sure it's -29 instead of -21? That number doesn't make sense to me.

OpenStudy (anonymous):

I'm sure. I copied and pasted it and I just double checked

OpenStudy (anonymous):

Thats really weird, I learned this last year but I may be doing it wrong.

OpenStudy (anonymous):

I'm sorry I couldn't really help

OpenStudy (anonymous):

It's alright

OpenStudy (freckles):

m(x) = 4x − 11 and n(x) = x − 10 to find m(n(x)) that is to find m(x-10) since n(x)=4x-11 you just replace all the x's in m(x)=4x-11 with (x-10)

OpenStudy (freckles):